# DREAM: Scalable Red Teaming for Text-to-Image Generative Systems via Distribution Modeling

### v1

@@ Abstract
Despite the integration of safety alignment and external filters, text-to-image (T2I) generative models are still susceptible to producing harmful content, such as sexual or violent imagery. This raises serious concerns about unintended exposure and potential misuse. Red teaming, which aims to proactively identify diverse prompts that can elicit unsafe outputs from the T2I system (including the core generative model as well as potential external safety filters and other processing components), is increasingly recognized as an essential method for assessing and improving safety before real-world deployment. Yet, existing automated red teaming approaches often treat prompt discovery as an isolated, prompt-level optimization task, which limits their scalability, diversity, and overall effectiveness. To bridge this gap, in this paper, we propose DREAM, a scalable red teaming framework to automatically uncover diverse problematic prompts from a given T2I system. Unlike most prior works that optimize prompts individually, DREAM directly models the probabilistic distribution of the target system’s problematic prompts, which enables explicit optimization over both effectiveness and diversity, and allows efficient large-scale sampling after training. To achieve this without direct access to representative training samples, we draw inspiration from energy-based models and reformulate the objective into simple and tractable objectives. We further introduce GC-SPSA, an efficient optimization algorithm that provide stable gradient estimates through the long and potentially non-differentiable T2I pipeline. During inference, we also propose a diversity-aware sampling strategy to enhance prompt variety. The effectiveness of DREAM is validated through extensive experiments, demonstrating that it surpasses 9 state-of-the-art baselines by a notable margin across a broad range of T2I models and safety filters in terms of prompt success rate and diversity. Additionally, DREAM successfully uncovers failure cases in 4 real-world commercial T2I systems and enables more robust safety fine-tuning that generalizes to unseen harmful prompts.
@@ 1 Introduction
Text-to-image (T2I) generative models [1, 2, 3, 4] are driving a new wave of visual content creation, reshaping our expectations of what machines are capable of. Trained on large-scale datasets [5], these models capture rich associations between language and imagery, allowing them to produce high-quality images with simple text inputs (known as prompts). Their ease of use and impressive flexibility have driven rapid adoption across creative arts, entertainment, and social media, particularly among younger users such as teenagers [6, 7, 8]. However, the same datasets that enable this versatility also inevitably contain unsafe content (e.g., sexually explicit material and violence) due to their massive scale and web-crawled nature [9, 10]. As a result, the models also acquire the ability to produce images of harmful content during real use, raising serious ethical, legal, and accountability concerns [11, 12].
@@
To mitigate these risks, a growing number of efforts from both academia and industry [13, 10, 14, 15, 16] have focused on improving the safety of T2I generative models. One popular approach is safety alignment, also referred to as unsafe concept erasure in the T2I literature [13, 10, 14, 17], which fine-tunes the model using a curated set of unsafe prompts or images to suppress undesirable generations. This process helps steer the model toward harmless outputs: for example, returning a clothed figure even when prompted with “a nude person”. In addition, commercial companies like Stability AI [18] and Ideogram [19] also employ proprietary safety filters (e.g., NSFW image detectors) to block generation attempts when unsafe content is detected. These filters, when combined with the core generative model and other processing components, constitute the deployed T2I system. However, while these techniques show promising results in controlled environments, they remain imperfect when applied in practice. For example, both real-world users and researchers [20, 21] have reported that prompts unseen during training (e.g., implicit references to sensitive content), or even totally benign inputs (e.g., “the origin of woman”), may still escape moderation and lead to unsafe outputs. These observations highlight the limitations of current methods under open-ended inputs and the urgent need for proactive mechanisms to expose safety vulnerabilities of T2I generative systems before real-world deployment.
@@
One emerging solution to proactively identify such blind spots is red teaming, where model owners (e.g., developers) simulates the behavior of real-world users to generate various testing prompts, aiming to systematically probe the model’s failure modes before deployment. In the context of T2I generative systems, red teaming typically attempts to find a diverse set of problematic prompts that can elicit unsafe or policy-violating outputs despite potential safeguards [20, 21]. By doing so, it not only serves as an evaluation tool for stress-testing the system’s safety and trustworthiness under open-ended inputs [20], but also provides valuable references for future improvement [22]. As a result, red teaming is increasingly recognized as a critical practice, with major companies like Google [23] initiating human-in-the-loop red teaming programs. At the same time, regulatory bodies are increasingly emphasizing rigorous safety testing before deployment, as reflected in the EU AI Act [24] and the U.S. NIST AI Risk Management Framework [25], alongside similar official efforts around the world [26, 27].
@@
While early red teaming practice relied on human experts, recent works [20, 21, 28] have shifted toward automated red teaming, aiming to discover problematic prompts without human oversight. For example, FLIRT [28] employs a large language model (LLM) to generate an initial prompt and iteratively rewrite it toward unsafe outputs, while P4D [20] starts with a moderated unsafe prompt and applies token-level gradient-based substitutions to penetrate safety alignment. However, these methods often struggle to balance the success rate with prompt diversity, and can be prohibitively slow and costly to scale (e.g., P4D [20] takes $\sim$30 minutes to optimize a single prompt). These limitations underscore the urgent need for a scalable red teaming method that can efficiently generate a large, diverse set of effective problematic prompts.
@@
In this paper, we present the first attempt towards bridging the aforementioned gap. Our method is driven by a unified insight into the shared limitations of previous works: they treat red teaming as a prompt-to-prompt discrete optimization problem, where each prompt is optimized independently yet without accumulating global knowledge across runs. Built upon this understanding, we propose Distributional Red tEAming via energy-based Modeling (DREAM), which directly models the probabilistic distribution of the target model’s unsafe prompts via training a parameterized prompt generator (e.g., an autoregressive LLM). In contrast to previous approaches, our formulation enables explicit optimization of both success and diversity, supports global updates to the modeled prompt distribution, and allows efficient large-scale sampling after training.
@@
However, modeling the target prompt distribution is challenging, as it is tightly coupled to the specific T2I system and lacks sufficiently representative samples, making direct training infeasible. To overcome this, we draw inspiration from energy-based models [29], and decompose the originally intractable training objective into two surprisingly simple ones that allow effective distribution learning without direct sample access. Moreover, to enable effective and efficient gradient-based optimization for these objectives under long and potentially non-differentiable pipelines, we introduce Gradient-Calibrated Simultaneous Perturbation Stochastic Approximation (GC-SPSA), an efficient zero-th order optimization method based on SPSA [30]. Specifically, it estimates gradients using only forward evaluations and further improves stability via a history-aware calibration mechanism. We also provide theoretical analysis and convergence guarantees to support the use of GC-SPSA for optimizing our objectives. Finally, we propose a novel adaptive temperature scaling strategy method to further increase coverage at inference time.
@@
We conduct extensive experiments on 5 state-of-the-art (SOTA) safety-aligned diffusion models, 4 safety filters, and compare DREAM with 9 SOTA baselines across two unsafe categories. The results show that DREAM consistently outperforms all baselines in terms of prompt success rate with a notable margin and approaches human-level diversity. DREAM also generalizes well to advanced T2I models (e.g., SDXL, SD v3) and 6 other NSFW themes, and remains effective even under strong combinational defenses or aggressive filters. Furthermore, an IRB-approved user study confirms that DREAM produces more effective and diverse prompts than prior methods. In addition, case studies demonstrate that DREAM can transfer to four commercial T2I platforms with unknown safety mechanisms. Finally, prompts generated by DREAM significantly enhance safety fine-tuning, enabling models to resist both seen and unseen harmful prompts better than those trained on other baselines.
@@
  • We revisit existing red teaming methods and identify a shared limitation: they treat prompt discovery as isolated, prompt-to-prompt optimization without global modeling, restricting their scalability and overall effectiveness.
  • We introduce DREAM, a scalable and distribution-aware red teaming framework that learns a probabilistic model over unsafe prompts using energy-based modeling. We further propose GC-SPSA, a novel zero-order optimization method that supports effective and efficient training, along with adaptive inference strategies for broader coverage. Theoretical analyses and global convergence guarantees are provided to support our framework.
  • We conduct comprehensive evaluations across 5 safety-aligned T2I models, 4 safety filters, and 9 SOTA baselines, demonstrating that DREAM achieves higher prompt success rate and matches human-level diversity. We also show DREAM ’s ability to expose failure cases in 4 commercial T2I platforms and to improve safety fine-tuning with strong generalization to unseen harmful prompts.
@@ 2.1 Text-to-Image Generative Models
Text-to-image (T2I) generative models have become a cornerstone of modern visual synthesis, enabling users to create highly detailed images from natural language descriptions. Among various generative paradigms, diffusion models [31, 3, 1, 2, 4] have emerged as the dominant approach due to their superior training stability, generation quality, and controllability. Diffusion models operate by iteratively denoise random noise into coherent images, often conditioned on texts, making them particularly effective for large-scale training and text-controlled generation. Building upon this, a wide range of open-source (e.g., Stable Diffusion family [4]) and commercial systems (e.g., DeepAI [32], DALL·E 3 [33], Midjourney [34], Ideogram [19]) have been developed, providing state-of-the-art generation experiences under user-friendly graphical interfaces.
@@ 2.2 Unsafe Generation & Mitigation
The success of modern T2I models relies heavily on large datasets. For instance, Stable Diffusion is trained on LAION-5B [5], a web-scraped set of over 5 billion image-text pairs, while commercial models like Ideogram use even larger private datasets [19]. These datasets support powerful multimodal learning but also contain harmful content such as harmful imagery and copyrighted materials [35, 36]. This can be absorbed and reproduced by these models, raising ethical and legal issues, especially as these tools become more accessible and popular among children and adolescents, who may suffer psychological harm, safety risks, and disrupted development from exposure [11, 9, 37].
@@
In response to these concerns, several mitigation strategies have emerged, which can be broadly categorized into two lines: model safety alignment and inference-time safety filtering. Model safety alignment [10, 14] refers to techniques that tune the diffusion model’s parameters directly to suppress its ability in producing unsafe content. This is typically achieved by collecting a curated set of harmful prompts or images and reinforcing the model to “unlearn” them through methods such as adversarial training [38], supervised fine-tuning [10, 13], or model editing [14, 17]. For example, CA [13] fine-tunes the diffusion model to match the image distribution of an unsafe target concept (e.g., “a nude woman”) to that of a safe anchor concept (e.g., “woman”). As a result, the model learns to resist prompts that are the same or similar to training-time target concepts and generates a safe image instead. In contrast, safety filters [16, 39, 15] act as external control mechanisms during inference. They can operate in prompt-level or image-level, aiming to detect and block unsafe content before or after generation. A representative case is the Safety Checker (SC) [15] employed in Stable Diffusion models, which compares the generated image with a set of predefined sensitive concepts and blocks outputs that exceed a similarity threshold.
@@
While these approaches have demonstrated effectiveness with acceptable trade-offs in benign performance under their own evaluation protocols, their robustness in real-world scenarios has been frequently challenged by a growing body of recent research and user reports. For instance, text-based safety filters can be bypassed using simple synonym substitutions (e.g., changing “gun” to “sidearm”) [40], while image-based filters may lose effectiveness under subtle alterations in image styles, compositions, or rendering (e.g., “a nude woman in colored painting”) [41]. Besides, safety-aligned models may perform well when the prompts contain explicit words (e.g., “nude” or “sexy”), but still fail to handle veiled expressions, metaphors or context-related implications such as “a woman looking seductive,” “a cute Japanese movie star,” [42, 43, 20], which are unseen during unlearning. In addition to these scattered findings, recent research [44, 45, 46, 47] has developed various optimization methods to transform a given rejected prompt into a minimally modified variant that bypasses safety mechanisms, a technique known as adversarial jailbreak attacks. These diverse failure patterns on different safety mechanisms suggest it is crucial to proactively test and improve the system’s safety before real-world deployment.
@@ 2.3 Red Teaming for Text-to-Image Models
The concept of “red teaming” originated during the Cold War era in the 1960s as a form of structured military system testings and has since expanded to fields like cybersecurity, airport security, software engineering, and recently to AI and ML systems [48]. For generative models, red teaming typically involves simulating real user behavior to explore the system and find prompts that produce harmful or policy-violating outputs [49, 50, 51]. Unlike jailbreak attacks that tweak known unsafe prompts to evasive variants [44, 46], red teaming focuses on broader exploration to reveal diverse or unexpected failure modes [21]. It is now a key part of responsible model development [23] and is increasingly emphasized in recent regulatory frameworks [24, 25, 26].
One predominant form of red teaming is manual construction. For example, the I2P dataset [44] was formed by collecting and filtering harmful prompts from various forums through a mix of automatic tools and human curation. Similarly, Google’s Adversarial Nibbler Challenge [23] invited participants to attack real-world T2I models and selected high-quality prompts based on their effectiveness and diversity. Commercial providers also employ in-house or external experts to manually test models for discovering failure modes [52]. While such methods can surface unexpected and model-specific vulnerabilities, they rely heavily on human effort and lack automation, making them inefficient and expensive to conduct.
To this end, several methods for automated red teaming have been proposed [20, 28, 21, 53]. These methods typically adopt paradigms and techniques similar to jailbreak attacks and transform a set of initial prompts to harmful ones, using methods like token-level substitution [20, 53] and LLM-rewrite [28, 21]. However, as we will identify in the following section, this inherited formulation will inherently limit their effectiveness, exploration space, and efficiency, making them suboptimal for scalable red teaming.

### v2

@@ Abstract
Despite the integration of safety alignment and external filters, text-to-image (T2I) generative systems are still susceptible to producing harmful content, such as sexual or violent imagery. This raises serious concerns about unintended exposure and potential misuse. Red teaming, which aims to proactively identify diverse prompts that can elicit unsafe outputs from the T2I system, is increasingly recognized as an essential method for assessing and improving safety before real-world deployment. However, existing automated red teaming approaches often treat prompt discovery as an isolated, prompt-level optimization task, which limits their scalability, diversity, and overall effectiveness. To bridge this gap, in this paper, we propose DREAM, a scalable red teaming framework to automatically uncover diverse problematic prompts from a given T2I system. Unlike prior work that optimizes prompts individually, DREAM directly models the probabilistic distribution of the target system’s problematic prompts, which enables explicit optimization over both effectiveness and diversity, and allows efficient large-scale sampling after training. To achieve this without direct access to representative training samples, we draw inspiration from energy-based models and reformulate the objective into a simple and tractable form. We further introduce GC-SPSA, an efficient optimization algorithm that provides stable gradient estimates through the long and potentially non-differentiable T2I pipeline. During inference, we also propose a diversity-aware sampling strategy to enhance prompt variety. The effectiveness of DREAM is validated through extensive experiments, demonstrating state-of-the-art performance across a wide range of T2I models and safety filters in terms of both prompt success rate and diversity. Our code is available at https://github.com/AntigoneRandy/DREAM.
@@ 1 Introduction
Text-to-image (T2I) generative models [croitoru2023diffusion, yang2023diffusion, rombach2022ldm, podell2024sdxl] are driving a new wave of visual content creation, reshaping our expectations of what machines are capable of. Trained on large-scale datasets [schuhmann2022laion], these models capture rich associations between language and imagery, allowing them to produce high-quality images with simple text inputs (known as prompts). Their ease of use and impressive flexibility have driven rapid adoption across creative arts, entertainment, and social media, particularly among younger users such as teenagers [wang2023factors, commonsense2024teens, ali2023constructing]. However, the same large-scale, web-crawled datasets that enable this versatility also inevitably contain not safe for work (NSFW) content (e.g., sexually explicit material) [schramowski2023sld, gandikota2023erasing]. As a result, the models also acquire the ability to produce harmful images during real use, raising serious ethical, legal, and accountability concerns [qu2023unsafe, ho2023development].
@@
To mitigate these risks, a growing number of efforts from both academia and industry [kumari2023ablating, gandikota2023erasing, gandikota2024unified, compvis2022sc, nsfw_text_classifier] have focused on improving the safety of T2I generative models. One popular approach is safety alignment, also referred to as unsafe concept erasure in the T2I literature [kumari2023ablating, gandikota2023erasing, gandikota2024unified, gong2024rece], which fine-tunes the model using a curated set of unsafe prompts or images to suppress undesirable generations. This process helps steer the model toward harmless outputs: for example, returning a clothed figure even when prompted with “a nude person”. In addition, commercial companies like Stability AI [stabilityai2025] and Ideogram [ideogram2025] also employ proprietary safety filters (e.g., NSFW image detectors) to block generation attempts when unsafe content is detected. These filters, when combined with the core generative model and other processing components, constitute the deployed T2I system. However, while these techniques show promising results in controlled environments, they remain imperfect when applied in practice. For example, both real-world users and researchers [chin2024p4d, li2024art] have reported that prompts unseen during training (e.g., implicit references to sensitive content), or even totally benign inputs (e.g., “the origin of woman”), may still escape moderation and lead to unsafe outputs. These observations highlight the limitations of current methods under open-ended inputs and the urgent need for proactive mechanisms to expose safety vulnerabilities of T2I generative systems before real-world deployment.
@@
One emerging solution to proactively identify such blind spots is red teaming, where model owners (e.g., developers) simulate the behavior of real-world users to generate various testing prompts, aiming to systematically probe the model’s failure modes before deployment. In the context of T2I generative systems, red teaming typically attempts to find a diverse set of problematic prompts that can elicit unsafe or policy-violating outputs despite potential safeguards [chin2024p4d, li2024art]. By doing so, it not only serves as an evaluation tool for stress-testing the system’s safety and trustworthiness under open-ended inputs [chin2024p4d], but also provides valuable references for future improvement [liu2024safetydpo]. As a result, red teaming is increasingly recognized as a critical practice, with major companies like Google [quaye2024advnibbler] initiating human-in-the-loop red teaming programs. At the same time, regulatory bodies are increasingly emphasizing rigorous safety testing before deployment, as reflected in the EU AI Act [euai2024] and the U.S. NIST AI Risk Management Framework [nist2023], alongside similar official efforts around the world [uk2024, singapore2024].
@@
While early red teaming relied on human experts, recent works [chin2024p4d, li2024art, mehrabi2024flirt] have shifted toward automated red teaming, aiming to discover problematic prompts with minimal human intervention. For example, FLIRT [mehrabi2024flirt] employs a large language model (LLM) to iteratively rewrite a seed prompt toward unsafe outputs, while P4D [chin2024p4d] starts with a moderated unsafe prompt and applies token-level gradient-based substitutions to penetrate safety alignment. However, these methods often struggle to balance the success rate with prompt diversity, and can be prohibitively slow and costly to scale. These limitations underscore the urgent need for a scalable red teaming method that can efficiently generate a large, diverse set of effective problematic prompts.
@@
In this paper, we present the first attempt to bridge the aforementioned gap. Our method is driven by a unified insight into the shared limitations of prior works: they treat red teaming as a prompt-to-prompt discrete optimization problem, where each seed prompt is optimized independently yet without accumulating global, distribution-level knowledge of the target model’s unsafe prompt distribution across runs. Built upon this understanding, we propose Distributional Red tEAming via energy-based Modeling (DREAM), which directly models the probabilistic distribution of the target model’s unsafe prompts by training a parameterized prompt generator (e.g., an autoregressive LLM). In contrast to previous approaches, our formulation enables explicit optimization of both success and diversity, supports global updates to the modeled prompt distribution, and allows efficient large-scale sampling after training.
@@
However, modeling the target prompt distribution is challenging, as it is tightly coupled to the specific T2I system and lacks sufficiently representative samples, making direct training infeasible. To overcome this, we draw inspiration from energy-based models [lecun2006tutorial] and decompose the originally intractable training objective into two simple sub-objectives that allow effective distribution learning without direct sample access. Moreover, to enable effective and efficient gradient-based optimization for these objectives under long and potentially non-differentiable pipelines, we introduce Gradient-Calibrated Simultaneous Perturbation Stochastic Approximation (GC-SPSA), an efficient zeroth-order optimization method based on SPSA [spall1992multivariate]. Specifically, it estimates gradients using only forward evaluations and further improves stability via a history-aware calibration mechanism. We also provide theoretical analysis and convergence guarantees to support GC-SPSA. Finally, we propose a novel adaptive temperature scaling method to further enhance coverage at inference time.
@@
We conduct extensive experiments on 5 state-of-the-art (SOTA) safety-aligned T2I models and 4 safety filters, comparing DREAM with 10 SOTA baselines across two unsafe categories. The results demonstrate that DREAM achieves superior performance in terms of prompt success rate while maintaining diversity comparable to human-written datasets. DREAM also generalizes well to advanced T2I models (e.g., SDXL, SD v3) and 6 other NSFW themes, and remains effective even under combinational defenses or aggressive filters. We further conduct a user study to confirm that the red-teaming prompts generated by DREAM are both diverse and highly effective from human perspectives. Moreover, case studies demonstrate that the prompts generated by DREAM transfer effectively to 4 commercial T2I platforms with unknown safety mechanisms. Finally, DREAM also enhances safety fine-tuning, enabling defended models to better resist both seen and unseen harmful prompts.
@@
  • We revisit existing red teaming methods and identify a shared limitation: they treat prompt discovery as an isolated, prompt-level optimization problem without global modeling, restricting their scalability and performance.
  • We introduce DREAM, a scalable and distribution-aware red teaming framework that learns a probabilistic model over unsafe prompts using energy-based modeling. We further propose GC-SPSA, a novel zeroth-order optimizer that supports effective and efficient training, along with adaptive inference strategies for broader coverage. We also provide theoretical analyses and global convergence guarantees to support our GC-SPSA framework.
  • We conduct comprehensive evaluations across 5 safety-aligned T2I models, 4 safety filters, and 10 SOTA baselines, showing that DREAM achieves superior prompt success rates while matching human-level diversity. We also show that DREAM can expose failure cases in 4 commercial T2I platforms and improve safety fine-tuning with notable generalization to unseen harmful prompts.
@@ 2.1 Text-to-Image Generative Models
Text-to-image (T2I) generative models have become a cornerstone of modern visual synthesis, enabling users to create highly detailed images from natural language descriptions. Among various generative paradigms, diffusion models [ho2020ddpm, rombach2022ldm, croitoru2023diffusion, yang2023diffusion, podell2024sdxl] have emerged as the dominant approach due to their superior training stability, generation quality, and controllability. Diffusion models operate by iteratively denoising random noise into coherent images, often conditioned on texts, making them particularly effective for large-scale training and text-controlled generation. Building upon this, a wide range of open-source (e.g., Stable Diffusion family [podell2024sdxl]) and commercial systems (e.g., DeepAI [deepai2025], DALL·E 3 [openai_dalle3_2023], Midjourney [midjourney2025], Ideogram [ideogram2025]) have been developed, providing state-of-the-art generation experiences through user-friendly graphical interfaces.
@@ 2.2 Unsafe Generation & Mitigation
The success of modern T2I models relies heavily on large datasets. For instance, Stable Diffusion is trained on LAION-5B [schuhmann2022laion], a web-scraped set of over 5 billion image-text pairs, while commercial models like Ideogram use even larger private datasets [ideogram2025]. These datasets support powerful multimodal learning but also contain harmful content such as sexual or violent imagery. Such content can be reproduced by these models, raising ethical and legal concerns. The risks become particularly significant as these tools grow more accessible and popular among children and adolescents, who may experience psychological harm, safety risks, and developmental disruptions due to exposure [qu2023unsafe, schramowski2023sld, guo2024moderating].
@@
In response to these concerns, several mitigation strategies have emerged, which can be broadly categorized into two lines: model safety alignment and inference-time safety filtering. Model safety alignment [gandikota2023erasing, gandikota2024unified] refers to techniques that tune the diffusion model’s parameters directly to suppress its ability to produce unsafe content. This is typically achieved by collecting a curated set of harmful prompts or images and reinforcing the model to “unlearn” them through methods such as adversarial training [zhang2024defensive], supervised fine-tuning [gandikota2023erasing, kumari2023ablating], or model editing [gandikota2024unified, gong2024rece]. For example, CA [kumari2023ablating] fine-tunes the diffusion model to match the image distribution of an unsafe target concept (e.g., “a nude person”) to that of a safe anchor concept (e.g., “person”). As a result, the model learns to resist prompts that are the same or similar to training-time target concepts and generates safe images instead. In contrast, safety filters [nsfw_text_classifier, chhabra_nsfw_detection_dl, compvis2022sc] act as external control mechanisms during inference. They can operate at prompt-level or image-level, aiming to detect and block unsafe content before or after generation. A representative case is the Safety Checker (SC) [compvis2022sc] employed in Stable Diffusion models, which compares the generated image with a set of predefined sensitive concepts and blocks outputs that exceed a similarity threshold.
@@
While these approaches have demonstrated effectiveness with acceptable trade-offs in benign performance under their own evaluation protocols, their robustness in real-world scenarios has been frequently challenged by a growing body of recent research and user reports. For instance, text-based safety filters can be bypassed using simple synonym substitutions [ba2024surrogateprompt], while image-based filters may lose effectiveness under subtle alterations in image styles, compositions, or rendering [reddit_dalle3_jailbreak]. Moreover, while safety-aligned models perform well when the prompts contain in-distribution explicit words, they may still fail to handle out-of-distribution veiled expressions, metaphors, or context-related implications [li2024safegen, meng2025concept, chin2024p4d], which are unseen during unlearning. In addition to these scattered findings, recent research [yang2024sneakyprompt, yang2024mma, huang2025perception, zhuang2023qfattack] has developed various optimization methods to transform a given rejected unsafe prompt into evasive variants to bypass safety mechanisms, a technique known as adversarial jailbreak attacks. These diverse failure patterns across different safety mechanisms suggest it is crucial to proactively test and improve the T2I generative system’s safety before real-world deployment.
@@ 2.3 Red Teaming for Text-to-Image Models
The concept of “red teaming” originated during the Cold War era in the 1960s as a form of structured military system testing and has since expanded to fields like cybersecurity, airport security, software engineering, and more recently to AI and ML systems [feffer2024red]. For generative models, red teaming typically involves simulating real user behavior to explore the system and find prompts that produce harmful or policy-violating outputs [verma2024operationalizing, hong2024curiosity, perez2022red]. Unlike jailbreak attacks that tweak known unsafe prompts into evasive variants [yang2024sneakyprompt, huang2025perception], red teaming focuses on broader exploration to reveal diverse or even unexpected failure modes [li2024art]. It is now a key part of responsible model development [quaye2024advnibbler] and is increasingly emphasized in recent regulatory frameworks [euai2024, nist2023, uk2024].
One predominant form of red teaming is manual construction. For example, the I2P dataset [yang2024sneakyprompt] was formed by collecting and filtering harmful prompts from various forums through a mix of automatic tools and human curation. Similarly, Google’s Adversarial Nibbler Challenge [quaye2024advnibbler] invited participants to attack real-world T2I models and selected high-quality prompts based on their effectiveness and diversity. Commercial providers also employ in-house or external experts to manually test models for discovering failure modes [openaiRedTeaming]. While such methods can surface unexpected and model-specific vulnerabilities, they rely heavily on human effort and lack automation, making them inefficient and expensive to conduct.
To this end, several methods for automated red teaming have been proposed [chin2024p4d, mehrabi2024flirt, li2024art, tsai2024ring]. These methods typically adopt paradigms and techniques similar to jailbreak attacks and transform a set of initial prompts into harmful ones, using methods like token-level substitution [chin2024p4d, tsai2024ring] and LLM-rewrite [mehrabi2024flirt, li2024art]. However, as we will identify in the following section, this inherited formulation inherently limits their effectiveness, exploration space, and efficiency, making them suboptimal for scalable red teaming.